OpenStudy (anonymous):

Evaluate the integral sin x {f′(cos x) − 2} dx from 0 to pi/2 when f(0)=1 and f(1)=5

OpenStudy (anonymous):

\[\int_{0}^{\frac{\pi}{2}} Sin(x)[f\prime(Cos(x))-2] \;\;dx =f(Cos(x))+2Cos[x] \|_{0}^{\frac{\pi}{2}} \] So, \[f(Cos(\frac{\pi}{2}))+2Cos[\frac{\pi}{2}]-(f(Cos(0))+2Cos[0])\] then \[f(0)+2(0)-(f(1)+2)\] Eventually you end up with this \[f(0)-f(1)-2= 1-5-2=-6\]

OpenStudy (anonymous):

Thanks for the help, but it's not the right answer. You did help get me on the right track. :)

OpenStudy (anonymous):

is the answer 2? thats what im getting using a u=cos x substitution.

OpenStudy (anonymous):

Answer is 2, sorry I forgot the negative

OpenStudy (anonymous):

yes. Can you show me how to do it using u sub?

OpenStudy (anonymous):

posting in a sec, scanner is scanning.

OpenStudy (anonymous):

when you do the u sub, make sure to remember to switch the limits of integration. thats what always gets me >.< lol

OpenStudy (anonymous):

THANKS! :)

OpenStudy (anonymous):

\[f((g))'=f'(g(x))g'(x)\] so \[\int f((g))'dx= \int f'(g(x))g'(x) dx\] \[g(x)= cos(x)\] so \[g(x)= -sin(x)\] I forgot the negative but everything else is fine. ^^

OpenStudy (anonymous):

Correction \[g(x)'=-sin(x)dx\]

OpenStudy (anonymous):

∫ π 2 0 Sin(x)[f′(Cos(x))−2]dx=f(Cos(x))+2Cos[x]∥ π 2 0 So, f(Cos(π 2 ))+2Cos[π 2 ]−(f(Cos(0))+2Cos[0]) then f(0)+2(0)−(f(1)+2) Eventually you end up with this f(0)−f(1)−2=1−5−2=−6

OpenStudy (aravindg):

luckey can u help mE??

OpenStudy (anonymous):

yeah., i'll try to now m leaving, have some work u mail ur prblm at lokeshsonee@gmail.com